# Hybrid estimation for a mixed fractional Black-Scholes model with random effects from discrete time observations

### v3

@@ 3.2 Nonparametric estimation of random effects distribution function
$\hat{F}_{m,n,N}(x)=\sum\limits_{j=1}^{m}\hat{F}_{N}\left(x_{j}\right)\mathcal{L}_{j}(x),\ \forall x\in[-1,1],$ (3.12)
where $\hat{F}_{N}(y)=\frac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\left\{\widehat{\phi}_{i}\leq y\right\}}$ is the empirical distribution function of the estimators $\widehat{\phi}_{i}$. The remainder of this subsection is devoted to studying the asymptotic properties of the estimator $\hat{F}_{m,n,N}$.
The following proposition sets forward the asymptotic bias, variance and mean squared error (in short MSE) of $\hat{F}_{m,n,N}$.
@@
    the asymptotic bias of $\hat{F}_{m,n,N}$ is given by
    $\lim\limits_{n,N\to\infty}\mathrm{Bias}\left(\hat{F}_{m,n,N}(x)\right)=\frac{\pi}{m^{2}}T_{m}(x)\mu(x)+O(m^{-2}),$ (3.13)
@@
    the asymptotic variance of $\hat{F}_{m,n,N}$ is given by
    $\lim\limits_{n,N\to\infty}Var\left(\hat{F}_{m,n,N}(x)\right)=N^{-1}\sigma_{F}^{2}(x)+O\left(N^{-1}m^{-1/2}\right),$ (3.14)
@@
    the asymptotic mean squared error of $\hat{F}_{m,n,N}$, denoted by $AMSE\left(\hat{F}_{m,n,N}(x)\right)$, is given by
    $\mathrm{AMSE}\left(\hat{F}_{m,n,N}(x)\right)=\dfrac{\pi}{m^{4}}C_{1}(x)+N^{-1}\sigma_{F}^{2}(x)+O(m^{-4})+O\left(N^{-1}m^{-1/2}\right),$ (3.15)
@@
Remark 2.
The asymptotically optimal $m$ for estimating the random effects distribution function $F$ with respect to $\mathrm{AMSE}$ is $m_{opt}=N^{-\frac{2}{7}}$ which leads to an optimal $\mathrm{AMSE}$ of order of $O(N^{-\frac{8}{7}})$. In comparison, the classical kernel distribution function estimator achieves an optimal convergence rate of $O(N^{-\frac{4}{5}})$ with respect to $\mathrm{AMSE}$. Thus, our proposed estimator exhibits a faster convergence rate ensures a better estimation accuracy.
The next proposition deals with the uniform strong consistency of $\hat{F}_{m,n,N}$.
@@
Let assumptions (A1)(A3)(A1)-(A3) be satisfied. Then, if $m=o\left(\frac{\sqrt{N}}{\log N}\right)$, the following uniform convergence holds
$\left\|\hat{F}_{m,n,N}-F\right\|\xrightarrow[]{\text{a.s.}}0,\quad\text{as}\quad n,N\to\infty.$
@@
Under assumptions (A1)(A3)(A1)-(A3), if $N^{\frac{1}{2}}m^{-2}\to 0$ as $N\to\infty$, then, for each fixed $x\in[-1,1]$, the estimator $\hat{F}_{m,n,N}$ is asymptotically normal, that is
$\sqrt{N}\left(\hat{F}_{m,n,N}(x)-F(x)\right)\xrightarrow{d}\mathcal{N}\left(0,\sigma_{F}^{2}\right),\quad\text{as}\quad n,N\to\infty,$
where $\sigma_{F}^{2}(x)=F(x)\left(1-F(x)\right)$.
In the next proposition, we show that, under certain conditions, the proposed Lagrange estimator satisfies the Chung-Smirnov property, which characterizes its extreme fluctuations about $F$ as the number of subjects $N$ increases. This property was investigated for the classical empirical distribution function by Chung (1949) and Smirnov (1944).
@@
Under assumptions (A1)(A3)(A1)-(A3), if $m=o\left(\left(\frac{\log\log(N)}{N}\right)^{\frac{1}{4}}\right)$ then $\hat{F}_{m,n,N}$ satisfies the Chung-Smirnov property, that is
$\underset{N\to\infty}{\lim\sup}\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{m,n,N}(x)-F(x)\right|\leq 1,\ \text{a.s}.$
@@ 4 Numerical simulations
Estimation methods are tested for $N\in\{100,250,500\}$ subjects and $n\in\{250,500,1000\}$ observations per subject. The simulated results involve the mean and standard deviation (S.dev) of the estimators $\hat{H},\ \hat{\gamma}^{2}$ and $\hat{\sigma}^{2}$, as well as the random effect estimator $\hat{\phi}_{i}$ and are presented in Table 1.
@@
$F$ $(N,n)$ $\hat{H}$ $\hat{\gamma}^{2}$ $\hat{\sigma}^{2}$ $\hat{\phi}_{i}$
Mean S.dev Mean S.dev Mean S.dev True Mean Mean($\hat{\phi}_{i}$) S.dev
@@
$\hat{F}_{h}(x):=\dfrac{1}{N}\sum\limits_{i=1}^{N}K\left(\dfrac{x-\hat{\phi}_{i}}{h}\right),\quad\forall\ x\in\mathbb{R},$ (4.1)
@@
Here, the Gaussian kernel is chosen. For the Lagrange estimator, the optimal value of $m$, denoted $m_{\text{opt}}$, is selected via 5-fold cross-validation by minimizing the integrated squared error (ISE) on a held-out validation set. The estimator $\hat{F}_{m,n,N}$ is then constructed using this optimal $m$. For each distribution function example from (1) to (4), we use the relevant transformation from those introduced in section 2.2 and we test both estimators under two sample size regimes $N=100$ subjects with $n=250$ observations per-subject and $N=250$ subjects with $n=500$. The plots of both estimators are presented in the following figures.
@@
Figure 1: Qualitative comparison between the proposed estimator $\hat{F}_{m,n,N}$ and the Kernel estimator $\hat{F}_{h}$ for the distribution $\mathcal{B}(2,2)$.
@@
Figure 2: Qualitative comparison between the proposed estimator $\hat{F}_{m,n,N}$ and the Kernel estimator $\hat{F}_{h}$ for the distribution $\mathcal{G}(2,1)$.
@@
Figure 3: Qualitative comparison between the proposed estimator $\hat{F}_{m,n,N}$ and the Kernel estimator $\hat{F}_{h}$ for the distribution $\mathcal{N}(0.5,0.25)$.
@@
Figure 4: Qualitative comparison between the proposed estimator $\hat{F}_{m,n,N}$ and the Kernel estimator $\hat{F}_{h}$ for the distribution $0.5*\mathcal{N}(-2,1)+0.5*\mathcal{N}(3,0.5)$.
In order to numerically compare the performance of $\hat{F}_{m,n,N}$ and $\hat{F}_{h}$, we compute the mean ISE of both estimators which is given by
$\mathrm{ISE}\left(\hat{F}\right):=\displaystyle\int_{-1}^{1}\left(\hat{F}(x)-F(x)\right)^{2}d\,x,$
@@
Table 2: Mean ISE of $\hat{F}_{m,n,N}$ and $\hat{F}_{h}$. The bold values indicate the smallest values of mean ISE.
@@
  • From Figures 1-4, we first observe that the accuracy of both estimators improves with a larger sample size. We also observe that the Lagrange estimator $\hat{F}_{m,n,N}$ captures local curvature better and consistently follows the general shape of the true distribution function more smoothly than the kernel estimator $\hat{F}_{h}$.
@@
  • By increasing $(N,n)$ from $(100,250)$ to $(500,1000)$ the standard deviations of $\hat{H}$, $\hat{\gamma}^{2}$ and $\hat{\sigma}^{2}$ decrease. This reflects higher precision and illustrates the theoretical asymptotic consistency of estimators.
  • Across all distributions, the standard deviation of $\hat{\phi}_{i}$ decreases as both the number of subjects $N$ and the number of observations per subject $n$ increase. This implies that the variance of $\hat{\phi}_{i}$ tends to zero as the sample size increases.
@@ 5 Empirical Application to Cryptocurrency Markets
Finally, the distribution of random effects is recovered, on the basis of $\hat{\phi}_{i},i=1,\ldots,n$, by the Lagrange estimator (3.12).
@@
Ticker $(n,N)$ $\hat{H}$ $\hat{\gamma}^{2}$ $\hat{\sigma}^{2}$ $\hat{\phi}_{\text{mean}}$ $\hat{\phi}_{\text{std}}$
@@
As both the number of trajectories $N$ and the number of discrete $n$ increase, global parameters $(\hat{H},\hat{\gamma}^{2},\hat{\sigma}^{2})$ tend to stabilize their values across all cryptocurrencies, thus, confirming the theoretical consistency of the two-step asymptotic framework we introduced in section 3. The estimated Hurst exponents cluster between 0.61 and 0.66 for most of assets, which indicates that cryptocurrency returns exhibit long-range dependence and match our theoretical assumption $H\in(\frac{1}{2},\frac{3}{4})$. Hence, we can deduce that the shocks in the returns are carried forward over medium horizons rather than disappearing instantly, which is exactly a feature that the fractional component of the model captures. The fractional variance parameter $\hat{\gamma}^{2}$ demonstrates a wide range of variability among the assets and is quite low for ETH, LINK, and AVAX and relatively high for MATIC and especially LTC, thus implying that some assets have strong memory. On the other hand, the short-term volatility parameter $\hat{\sigma}^{2}$ is higher for MATIC and LINK and lower for both LTC and ETH. Moreover, the standard deviation of the random effects, $\hat{\phi}_{std}$, gets considerably smaller as we use more data. This very important decrease shows that with a larger sample size, our estimation method distinguishes better real, lasting differences in the trend of an asset from random noise. In small samples, random effects reflect not only actual differences across trajectories but also estimation noise. In contrast, in larger samples, the noise component gets smaller and the remaining dispersion corresponds to genuine, persistent asset-specific drifts.
@@ 6 Conclusion
In this work, we investigated a hybrid estimation framework for a mixed fractional Black-Scholes model incorporating both fixed global parameters and subject-specific random effects. By using moment-based statistics and discrete observations, we developed consistent and asymptotically normal estimators for model fixed parameters and random effects. We then addressed the estimation of the random effects distribution, particularly in cases where its support is compact, and we introduced a nonparametric approach based on Lagrange interpolation at Chebyshev-Gauss nodes. The estimator was shown to perform reliably and consistently as both the number of trajectories $N$ and the number of discrete observations per-subject increase. The robustness of our procedure was rigorously validated through a simulation study and an empirical application to cryptocurrency markets, demonstrating its practical utility in decomposing complex financial time series into persistent noise, transient volatility, and asset-specific fundamental trends.

### v4

@@ 3.2 Nonparametric estimation of random effects distribution function
$\widehat{F}_{m,n,N}(x)=\sum\limits_{j=1}^{m}\widehat{F}_{N}\left(x_{j}\right)\mathcal{L}_{j}(x),\ \forall x\in[-1,1],$ (3.15)
where $\widehat{F}_{N}(y)=\frac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\left\{\widehat{\phi}_{i}\leq y\right\}}$ is the empirical distribution function of the estimators $\widehat{\phi}_{i}$. The remainder of this subsection is devoted to studying the asymptotic properties of the estimator $\widehat{F}_{m,n,N}$.
The following proposition sets forward the asymptotic bias, variance and mean squared error (in short MSE) of $\widehat{F}_{m,n,N}$.
@@
    the asymptotic bias of $\widehat{F}_{m,n,N}$ is given by
    $\lim\limits_{n,N\to\infty}\mathrm{Bias}\left(\widehat{F}_{m,n,N}(x)\right)=\frac{\pi}{m^{2}}T_{m}(x)\mu(x)+O(m^{-2}),$ (3.16)
@@
    the asymptotic variance of $\widehat{F}_{m,n,N}$ is given by
    $\lim\limits_{n,N\to\infty}Var\left(\widehat{F}_{m,n,N}(x)\right)=N^{-1}\sigma_{F}^{2}(x)+O\left(N^{-1}m^{-\frac{1}{2}}\right),$ (3.17)
@@
    the asymptotic mean squared error of $\widehat{F}_{m,n,N}$, denoted by $AMSE\left(\widehat{F}_{m,n,N}(x)\right)$, is given by
    $\mathrm{AMSE}\left(\widehat{F}_{m,n,N}(x)\right)=\dfrac{\pi}{m^{4}}C_{1}(x)+N^{-1}\sigma_{F}^{2}(x)+O(m^{-4})+O\left(N^{-1}m^{-\frac{1}{2}}\right),$ (3.18)
@@
The next proposition deals with the uniform strong consistency of $\widehat{F}_{m,n,N}$.
@@
Let assumptions (A1)(A3)(A1)-(A3) hold and $m=m_{N}$ satisfy $m_{N}\to\infty$ and $\frac{\log m_{N}}{m_{N}^{2}}\to 0$ as $N\to\infty$. Then, Under the sequential regime $n\to\infty$ first, then $N\to\infty$, there exists a sequence $n=n_{N}\to\infty$ such that
$\bigl\|\widehat{F}_{m,n,N}-F\bigr\|\overset{a.s}{\longrightarrow}0,\qquad N\to\infty.$
@@
Assume (A1)(A3)(A1)-(A3). Let $m=m_{N}$ satisfy $N^{-\frac{1}{2}}\,\frac{\log m}{m^{2}}\;\longrightarrow\;0\quad\text{as }N\to\infty.$ Then, under the sequential asymptotic regime $n\to\infty$ followed by $N\to\infty$, for every fixed $x\in[-1,1]$,
$\sqrt{N}\bigl(\widehat{F}_{m,n,N}(x)-F(x)\bigr)\;\xrightarrow{d}\;\mathcal{N}\!\left(0,\sigma_{F}^{2}(x)\right),\qquad\sigma_{F}^{2}(x)=F(x)\bigl(1-F(x)\bigr).$
In the following theorem, we establish an explicit non-asymptotic upper bound for the mean squared error of $\hat{F}_{m,n,N}$, quantifying the contribution of the interpolation error, the finite-sample effect, and the error induced by estimating the random effects.
Theorem 4.
Assume (A1)-(A3) are satisfied and let $m=m_{N}$ and $n=n_{N}$. Let $\varepsilon_{N}\to 0$ be a sequence satisfying
$\max\limits_{1\leq i\leq N}|\widehat{\phi}_{i}-\phi_{i}|\leq\varepsilon_{N},\quad\text{a.s.}$
Then
$\mathbb{E}\left[\|\widehat{F}_{m,n,N}-F\|^{2}\right]\leq\frac{C_{1}(\log m_{N})^{2}}{m_{N}^{4}}+\frac{C_{2}}{N}+C_{3}\varepsilon_{N}^{2}+o\left(\frac{1}{m_{N}^{4}}+\frac{\varepsilon_{N}}{N}\right),$ (3.19)
where $C_{1}$, $C_{2}$, $C_{3}$ are positive constants depending on $\|f\|$, $\|f^{\prime}\|$ and $h$.
Remark 4.
A natural way to balance the contributions of the non-asymptotic bound terms is to choose $m_{N}\sim N^{\frac{1}{4}}$, yielding $\frac{(\log m_{N})^{2}}{m_{N}^{4}}\asymp\frac{(\log N)^{2}}{N}$. Moreover, under the sequential regime $n\to\infty$ followed by $N\to\infty$, Proposition 5 and Theorem 1 ensure that
$\hat{\phi}_{i}-\phi_{i}=O_{\mathbb{P}}\big(n_{N}^{H-1}\big)+O_{\mathbb{P}}\big(N^{-\frac{1}{2}}\big).$
Then, by choosing $n_{N}=N^{2}$ and under the condition $H<\frac{3}{4}$, we obtain $\varepsilon_{N}=O_{\mathbb{P}}\big(N^{-\frac{1}{2}}\big)$, so that $C_{3}\varepsilon_{N}^{2}\asymp N^{-1}$. With these choices, the three leading terms are all of order $N^{-1}$. Consequently, there exists a constant $C>0$ such that, for $N$ large enough,
$\mathbb{E}\big\|\hat{F}_{m_{N},n_{N},N}-F\big\|^{2}\;\leq\;C\,\frac{(\log N)^{2}}{N}.$
Hence, our non-asymptotic bound demonstrates that the estimator $\hat{F}_{m_{N},n_{N},N}$, which is based on plug-in estimators of the random effects, achieves, up to a logarithmic factor, the same convergence rate as the empirical c.d.f $\bar{F}_{N}$ would if the random effects were directly observed.
In the next proposition, we show that, under certain conditions, the proposed Lagrange estimator satisfies the Chung-Smirnov property, which characterizes its extreme fluctuations about $F$ as the number of subjects $N$ increases. This property was investigated for the empirical c.d.f by Chung (1949) and Smirnov (1944).
@@
Let assumptions (A1)(A1)-$(A3)$ hold. Assume that $n$ and $N$ tend to infinity sequentially and let $m=m_{N}\to\infty$ be such that $\left(\frac{2N}{\log\log N}\right)^{\frac{1}{2}}\frac{\log m}{m^{2}}\longrightarrow 0.$ Then $\widehat{F}_{m,n,N}$ satisfies the Chung-Smirnov property, that is
$\underset{N\to\infty}{\lim\sup}\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\widehat{F}_{m,n,N}(x)-F(x)\right|\leq 1,\ \text{a.s}.$
@@ 4 Numerical simulations
Estimation methods are tested for $N\in\{100,250,500\}$ subjects and $n\in\{250,500,1000\}$ observations per subject. The simulated results involve the mean and standard deviation (S.dev) of the estimators $\widehat{H},\ \widehat{\gamma}^{2}$ and $\widehat{\sigma}^{2}$, as well as the random effect estimator $\widehat{\phi}_{i}$ and are presented in Table 1.
@@
$F$ $(N,n)$ $\widehat{H}$ $\widehat{\gamma}^{2}$ $\widehat{\sigma}^{2}$ $\widehat{\phi}_{i}$
Mean S.dev Mean S.dev Mean S.dev True Mean Mean($\widehat{\phi}_{i}$) S.dev
@@
$\widehat{F}_{h}(x):=\dfrac{1}{N}\sum\limits_{i=1}^{N}K\left(\dfrac{x-\widehat{\phi}_{i}}{h}\right),\quad\forall\ x\in\mathbb{R},$ (4.1)
@@
Here, the Gaussian kernel is chosen. For each distribution function example from (1) to (4), we use the relevant transformation introduced in Section 2.2 and test both estimators under two sample size regimes: $N=100$ subjects with $n=250$ observations per subject and $N=250$ subjects with $n=500$ observations. For the choice of the interpolation ordermm, we use $K$-fold cross-validation. We first partition the estimated random effects $\{\widehat{\phi}_{i},i=1,\ldots,N\}$ into $K$ folds of approximately equal size, with index sets $I_{1},\ldots,I_{K}$. For each candidate $m\in\mathcal{M}$, and for each fold $k$:
  • we construct the Lagrange estimator $\widehat{F}^{(-k)}_{m}$ using all data points with indices in $\{1,\ldots,N\}\setminus I_{k}$.
  • we evaluate its prediction error on the validation fold via
    $\mathrm{VE}_{k}(m)=\frac{1}{|I_{k}|}\sum_{i\in I_{k}}\bigl[\widehat{F}^{(-k)}_{m}(\widehat{\phi}_{i})-\widehat{F}^{(k)}(\widehat{\phi}_{i})\bigr]^{2},$
    where $\widehat{F}^{(k)}$ denotes the empirical c.d.f. based only on observations in $I_{k}$.
The cross-validation score for $m$ is then
$\mathrm{VE}(m)=K^{-1}\sum_{k=1}^{K}\mathrm{VE}_{k}(m),$
and we select
$m_{\mathrm{CV}}=\arg\min_{m\in\mathcal{M}}\mathrm{VE}(m).$
In the presented simulations, we take $K=5$ and $\mathcal{M}=\{5,\ldots,20\}$, and we denote the selected value by $m_{\text{opt}}$. The final Lagrange estimator $\widehat{F}_{m,n,N}$ is then constructed using $m=m_{\text{opt}}$.
The plots of both estimators are presented in the figures that follow.
@@
Figure 1: Qualitative comparison between the proposed estimator $\widehat{F}_{m,n,N}$ and the Kernel estimator $\widehat{F}_{h}$ for the distribution $\mathcal{B}(2,2)$.
@@
Figure 2: Qualitative comparison between the proposed estimator $\widehat{F}_{m,n,N}$ and the Kernel estimator $\widehat{F}_{h}$ for the distribution $\mathcal{G}(2,1)$.
@@
Figure 3: Qualitative comparison between the proposed estimator $\widehat{F}_{m,n,N}$ and the Kernel estimator $\widehat{F}_{h}$ for the distribution $\mathcal{N}(0.5,0.25)$.
@@
Figure 4: Qualitative comparison between the proposed estimator $\widehat{F}_{m,n,N}$ and the Kernel estimator $\widehat{F}_{h}$ for the distribution $0.5*\mathcal{N}(-2,1)+0.5*\mathcal{N}(3,0.5)$.
In order to numerically compare the performance of $\widehat{F}_{m,n,N}$ and $\widehat{F}_{h}$, we compute the mean ISE of both estimators which is given by
$\mathrm{ISE}\left(\widehat{F}\right):=\displaystyle\int_{-1}^{1}\left(\widehat{F}(x)-F(x)\right)^{2}d\,x,$
@@
Table 2: Mean ISE of $\widehat{F}_{m,n,N}$ and $\widehat{F}_{h}$. The bold values indicate the smallest values of mean ISE.
@@
  • From Figures 1-4, we first observe that the accuracy of both estimators improves with a larger sample size. We also observe that the Lagrange estimator $\widehat{F}_{m,n,N}$ captures local curvature better and consistently follows the general shape of the true distribution function more smoothly than the kernel estimator $\widehat{F}_{h}$.
@@
  • By increasing $(N,n)$ from $(100,250)$ to $(500,1000)$ the standard deviations of $\widehat{H}$, $\widehat{\gamma}^{2}$ and $\widehat{\sigma}^{2}$ decrease. This reflects higher precision and illustrates the theoretical asymptotic consistency of estimators.
  • Across all distributions, the standard deviation of $\widehat{\phi}_{i}$ decreases as both the number of subjects $N$ and the number of observations per subject $n$ increase. This implies that the variance of $\widehat{\phi}_{i}$ tends to zero as the sample size increases.
@@ 5 Empirical Application to Cryptocurrency Markets
Finally, the distribution of random effects is recovered, on the basis of $\widehat{\phi}_{i},i=1,\ldots,N$, by the Lagrange estimator (3.15).
@@
Ticker $(n,N)$ $\widehat{H}$ $\widehat{\gamma}^{2}$ $\widehat{\sigma}^{2}$ $\widehat{\phi}_{\text{mean}}$ $\widehat{\phi}_{\text{std}}$
@@
As both the number of trajectories $N$ and the number of discrete $n$ increase, global parameters $(\widehat{H},\widehat{\gamma}^{2},\widehat{\sigma}^{2})$ tend to stabilize their values across all cryptocurrencies, thus, confirming the theoretical consistency of the two-step asymptotic framework we introduced in section 3. The estimated Hurst exponents cluster between 0.61 and 0.66 for most of assets, which indicates that cryptocurrency returns exhibit long-range dependence and match our theoretical assumption $H\in(\frac{1}{2},\frac{3}{4})$. Hence, we can deduce that the shocks in the returns are carried forward over medium horizons rather than disappearing instantly, which is exactly a feature that the fractional component of the model captures. The fractional variance parameter $\widehat{\gamma}^{2}$ demonstrates a wide range of variability among the assets and is quite low for ETH, LINK, and AVAX and relatively high for MATIC and especially LTC, thus implying that some assets have strong memory. On the other hand, the short-term volatility parameter $\widehat{\sigma}^{2}$ is higher for MATIC and LINK and lower for both LTC and ETH. Moreover, the standard deviation of the random effects, $\widehat{\phi}_{std}$, gets considerably smaller as we use more data. This very important decrease shows that with a larger sample size, our estimation method distinguishes better real, lasting differences in the trend of an asset from random noise. In small samples, random effects reflect not only actual differences across trajectories but also estimation noise. In contrast, in larger samples, the noise component gets smaller and the remaining dispersion corresponds to genuine, persistent asset-specific drifts.
@@
In this work, we investigated a hybrid estimation framework for a mixed fractional Black-Scholes model incorporating both fixed global parameters and subject-specific random effects. By using moment-based statistics and discrete observations, we developed consistent and asymptotically normal estimators for model fixed parameters and random effects. We then addressed the estimation of the random effects distribution, particularly in cases where its support is compact, and we introduced a nonparametric approach based on Lagrange interpolation at Chebyshev-Gauss nodes. The estimator was shown to perform reliably and consistently as both the number of trajectories $N$ and the number of discrete observations per-subject increase. The robustness of our procedure was rigorously validated through a simulation study and an empirical application to cryptocurrency markets, demonstrating its practical utility in decomposing complex financial time series into persistent noise, transient volatility, and asset-specific fundamental trends. Extensions of the present work could address the more challenging case $H\geq\frac{3}{4}$, where the asymptotic limits become non‑Gaussian, requiring different normalizations and inference techniques.